# Optimal Modified Feedback Strategies in LQ Games under Control Imperfections

### v3

@@ Abstract
Game-theoretic approaches and Nash equilibrium have been widely applied across various engineering domains. However, practical challenges such as disturbances, delays, and actuator limitations can hinder the precise execution of Nash equilibrium strategies. This work investigates the impact of such implementation imperfections on game trajectories and players’ costs in the context of a two-player finite-horizon linear quadratic (LQ) nonzero-sum game. Specifically, we analyze how small deviations by one player, measured or estimated at each stage, affect the state and cost function of the other player. To mitigate these effects, we propose an adjusted control policy that optimally compensates for the deviations under the stated information structure and can, under certain conditions, exploit them to improve performance. Rigorous mathematical analysis and proofs are provided, and the effectiveness of the proposed method is demonstrated through a representative numerical example.
@@ I Introduction
Game theory is a powerful analytical framework for modeling rational decision-making in complex multi-agent systems, particularly when interactions are strategic and agents’ decisions influence each other [shamma2020game, basar1999dynamic]. Such interactions often involve conflicting objectives, motivating the development of solution concepts to characterize equilibrium behavior. Among these, the Nash equilibrium is a well-established notion, defined as a set of strategies from which no participant can unilaterally deviate to achieve a better outcome [nash1950equilibrium].
@@
Thanks to its solid theoretical foundation, Nash equilibrium has been widely applied in diverse engineering domains, including robotics [chiu2021encoding], traffic management [Oszczypała2023Nash], and sensor networks [Zheng2023game], as well as in emerging areas such as energy management [YANG2020216nashQ, abdolmohammadi2024sizing], autonomous vehicle navigation [nazari2024powertrain, bouzidi2025interaction, abtahi2025multi], and human–robot interaction [soltanian2025peerawarecostestimationnonlinear].
@@
Despite its wide adoption, implementing Nash strategies in real-world engineering systems faces significant challenges due to uncertainties, disturbances, delays, and actuator imperfections [van-den=broek2023, AMATO2002507_guaranteeing, ma2024ϵ]. Extensions such as the $\varepsilon$-equilibrium relax optimality by allowing bounded deviations from Nash strategies [TANAKA1991413], enabling numerical approximations [NORTMANN20231760, alvarez2009urban] and supporting further studies on sensitivity and robustness [van-den=broek2023, AMATO2002507_guaranteeing, Jimenez01072006, tan2005sensitivity, hernandez2024ϵ].
@@
Robustness of Nash equilibrium to uncertainty has been studied from various perspectives. For example, model uncertainties in differential games were addressed in [van-den=broek2023, AMATO2002507_guaranteeing, hernandez2024ϵ], where [van-den=broek2023] designed a finite-horizon feedback Nash strategy for Linear Quadratic (LQ) games, [AMATO2002507_guaranteeing] extended the scope to infinite-horizon settings, and [hernandez2024ϵ] relaxed the LQ structure via weaker convexity assumptions. Other works addressed different uncertainty sources: [Jimenez01072006] proposed adaptive control to mitigate measurement noise, and [tan2005sensitivity] analyzed the effect of cost parameter variations in discrete-time games.
@@
Nonetheless, the sensitivity of Nash equilibrium to player-specific control deviations—arising from implementation imperfections—remains largely unexplored. Such deviations, e.g., due to actuator errors or delays, are common in engineering systems and can directly degrade performance. Moreover, the discrete-time setting has received little attention in this context. This work analytically investigates how such deviations influence system trajectories and player costs.
@@
To address these effects, we propose a modified feedback policy that both mitigates performance loss and exploits available deviation information. While prior adaptive multi-agent methods exist [guerrero2021openloop, wang2022risk], they do not explicitly target compensation for execution-time control deviations of a specific player, which is addressed in this work.
@@
  • An optimal modified feedback control policy that compensates for such deviations and ensures performance that is never worse—and often strictly better—than the original feedback Nash equilibrium.
  • A complete theoretical analysis, supported by a representative numerical example, that highlights the robustness of Nash equilibrium under execution imperfections.
@@
Notation. $\mathbb{R}$ and $\mathbb{Z}$ denote the sets of real and integer numbers, respectively. For a matrix $A$, $A^{\top}$ denotes its transpose; for a square matrix $B$, $B\succ 0$ ($B\succeq 0$) means $B$ is positive definite (positive semi-definite), and $B^{-1}$ denotes its inverse. For $i\in N_{p}=\{1,2\}$, the subscript -i refers to $N_{p}\setminus\{i\}$, i.e., all indices except $i$. The superscript denotes the exact optimal solution.
@@ II Preliminaries
We focus on finite-horizon linear quadratic (LQ) games [Engwerda2009lq] with discrete-time dynamics
@@
We use Nash equilibrium to mean a strategy profile $U_{i}$, where no player can reduce its cost by a unilateral deviation, i.e. $J_{i}(x_{0},U_{i}^{\star},U_{-i}^{\star})\leq J_{i}(x_{0},U_{i},U_{-i}^{\star})$ [nash1950equilibrium].
@@
The FNE gains $\{K^{\star}_{i,k}\}$ are obtained by a backward Riccati recursion coupled across players [salizzoni2025bridging]. Let $\tilde{P}_{i,N}=Q_{i,N}$ and define, for each $k\in\{0,\dots,N-1\}$ (computed backward),
@@
Under the standard assumptions $Q_{i},Q_{i,N}\succeq 0$, $R_{i}\succ 0$ and nonsingularity of the gain system in Definition 1 at each $k$, the backward sweep $k=N-1\to 0$ yields a unique FNE [Engwerda2009lq, Sec. 6].
@@
Under conditions 1-3, a unique finite-horizon Nash equilibrium exists. We assume these conditions henceforth [LAA_Jungers_2013].
@@ III First-Order Sensitivity of Player 1 Cost under Opponent Deviations
as Definition 1. Substituting (5) into the dynamics (1) yields the time-varying closed-loop matrix
@@
with the convention $\Phi(j,j):=I$. Then,
@@ IV Control Policy Design with Disturbance Compensation
In practice, the control implementation is rarely exact. Differences between the commanded and implemented inputs typically arise from (i) actuator/servo dynamics and inner-loop tracking (first-order lags, rate/slew limits), (ii) computation/communication latency and sample-and-hold effects, and (iii) command filtering/smoothing or safety layers that reshape inputs. This creates a measured input disturbance $w_{k}\;:=\;B_{2}\big(u_{2,k}-u_{2,k}^{\star}\big)$ in exactly the channel exposed by our sensitivity result.
We, therefore, model $w_{k}$ with simple, identifiable dynamics. Then, we equip Player 1 with a Compensated Feedback (CF) law, which includes a feedforward term $L_{k}w_{k}$ to cancel the predictable component of the disturbance at the actuator, so the feedback term $K_{k}$ regulates the residuals instead of reacting after the state has already moved. We first treat a physically grounded actuator-lag model, then narrow it down to a specific case of slow-varying error dynamics.
@@ IV-A Compensated Feedback under Lagged Execution Error Dynamics
Inspired by the Disturbance-Based State Compensation (DBSC) approach [rugh1996linear], which is commonly used in traditional state-feedback control design, and given Assumption 1, we define the augmented state as $z_{k}:=[x_{k}^{\top}\ w_{k}^{\top}]^{\top}$.
@@
By (21) in Theorem 1, the gains $\{K_{k},L_{k}\}$ are optimal over the causal affine class (16), which includes the REF baseline $u_{1,k}^{\text{REF}}$ as the special case $L_{k}\equiv 0$. Consequently, execution-time deviations by Player 2 create an opportunity for Player 1 to outperform its own equilibrium-derived feedback (REF) by adopting the compensated law. This does not contradict the definition of a Nash equilibrium: the FNE is defined for the nominal game without implementation deviations, whereas here Player 2’s realized input differs from its Nash command. The approach is practically relevant because such deviations routinely arise from disturbances, physical limits, and execution imperfections.
@@ IV-B Specialization: Slowly-Varying Error Dynamics
which generally differ from the FNE dynamics $x_{k+1}=\big(A-B_{1}K^{\star}_{1,k}-B_{2}K^{\star}_{2,k}\big)\,x_{k}$. Stability and performance must therefore be assessed for the Linear Time Invariant (LTV) matrix sequence produced by the augmented Riccati recursion.
@@ V Numerical Example: Two Carts Coupled by a Spring–Damper
Let $x:=\begin{bmatrix}p_{1}&v_{1}&p_{2}&v_{2}\end{bmatrix}^{\top}$ in which $p_{1}$ and $p_{2}$ are the positions and $v_{1}$ and $v_{2}$ are the velocities of Player 1 and 2, respectively. The continuous-time system dynamics are $\dot{x}=A_{c}x+B_{1,c}u_{1}+B_{2,c}u_{2}$ with the matrices:
@@
We compare three implementations over the same horizon and initial condition. Case I (FNE): both players apply their finite-horizon feedback Nash laws computed from the coupled Riccati recursion; no actuator lag is present, and Player 2’s intended command is applied directly. Case II (REF): Player 2’s intended command is subjected to a first-order discrete lag with factor $\alpha$, but Player 1 continues to use its nominal feedback policy; this isolates the performance loss due to the execution lag. Case III (CF): the same actuator lag is present on Player 2, while Player 1 designs a compensated feedback by solving a finite-horizon Riccati recursion on the exact augmented realization of the lagged plant, yielding time-varying gains that act on the state and on the lag-induced mismatch signal.
@@
Under the compensated feedback (Case III, dotted), Player 1 modifies its input relative to the nominal FNE action to proactively counteract the actuation shortfall of Player 2. The resulting trajectories stay appreciably closer to the FNE benchmark—especially for Player 1—confirming that CF recovers a substantial fraction of the lag-induced performance loss. Player 2’s cost, however, rises further because it neither alters its strategy nor removes the actuator lag, highlighting the asymmetric benefit in a general-sum game.
Numerically (Table II), Player 1’s cost increases by $51.36\%$ when lag is present without compensation (Case II vs. Case I). With CF (Case III), this penalty is reduced: $J_{1}$ drops from $492.40$ to $439.67$, a $10.71\%$ improvement relative to the uncompensated case, recovering about $31.56\%$ of the loss from the nominal baseline.
@@
Table II: Finite-horizon costs under the three cases ($\tau=0.8$ s, $\alpha\approx 0.9814$). CF (Case III) reduces Player 1’s lag penalty relative to Case II.
@@ VI Conclusion
In this paper, we analyzed the impact of control implementation errors in finite-horizon, two-player, nonzero-sum LQ games. Our analysis focused on how deviations from the Nash equilibrium influence system stability and performance. We quantified the deviation of the system trajectory from the Nash trajectory and assessed the increase in the cost function when one player fails to follow the exact equilibrium strategy. Our analysis demonstrated that, when a Nash equilibrium exists, the system trajectory remains close to the Nash trajectory for small deviations from the equilibrium strategy.
Additionally, we proposed an adjusted control policy to mitigate the adverse effects of deviations by the other player from the Nash policy. This adjustment, grounded in the principle of dynamic programming, offers superior performance compared to strictly adhering to the pure Nash strategy. By optimally modifying the player’s response, the proposed policy minimizes the negative impact of the other player’s deviations and capitalizes on these deviations to achieve enhanced outcomes. The effectiveness of the proposed approach was validated through a known example. These findings highlight the potential for improved performance in real-world applications where executing the exact Nash strategy is impractical due to practical challenges such as actuator limitations or external disturbances.
@@
A promising avenue for future research is the development of online estimation techniques combined with adaptive control laws that can anticipate deviations of the other player from the Nash strategy and dynamically adjust to these errors in real time. Furthermore, examining the effects of more complex deviation patterns—such as time-varying, correlated, or structured errors—could offer deeper insights into the robustness and adaptability of game-theoretic control strategies.

### v4

@@ Abstract
Game-theoretic approaches and Nash equilibrium have been widely applied across various engineering domains. However, practical challenges such as disturbances, delays, and actuator limitations can hinder the precise execution of Nash equilibrium strategies. This work investigates the impact of such implementation imperfections on game trajectories and players’ costs in the context of a two–player finite–horizon linear quadratic (LQ) nonzero-sum game. Specifically, we analyze how small deviations by one player, measured or estimated at each stage affect the state trajectory and the other player’s cost. To mitigate these effects, we construct a compensation law for the influenced player by augmenting the nominal game with the measurable deviation dynamics. The resulting policy is shown to be optimal within a causal affine policy class, and, for sufficiently small deviations, it locally outperforms the uncompensated equilibrium-derived feedback. Rigorous analysis and proofs are provided, and the effectiveness of the proposed approach is demonstrated through a representative numerical example.
@@ I Introduction
Game theory is a powerful analytical framework for modeling rational decision-making in complex multi-agent systems, particularly when interactions are strategic and agents’ decisions influence each other [17, 5]. Such interactions often involve conflicting objectives, motivating the development of solution concepts to characterize equilibrium behavior. Among these, the Nash equilibrium is a well-established notion, defined as a set of strategies from which no participant can unilaterally deviate to achieve a better outcome [12].
@@
Thanks to its solid theoretical foundation, Nash equilibrium has been widely applied in diverse engineering domains, including robotics [7], traffic management [Oszczypała2023Nash], and sensor networks [24], as well as in emerging areas such as energy management [23, 1], autonomous vehicle navigation [13, 6, 2], and human–robot interaction [18].
@@
Despite its wide adoption, implementing Nash strategies in real-world engineering systems faces significant challenges due to uncertainties, disturbances, delays, and actuator imperfections [21, 4, ma2024ϵ]. Extensions such as the $\varepsilon$-equilibrium relax optimality by allowing bounded deviations from Nash strategies [20], enabling numerical approximations [14, 3] and supporting further studies on sensitivity and robustness [21, 4, 10, 19, hernandez2024ϵ].
@@
Robustness of Nash equilibrium to uncertainty has been studied from various perspectives. For example, model uncertainties in differential games were addressed in [21, 4, hernandez2024ϵ], where [21] designed a finite-horizon feedback Nash strategy for Linear Quadratic (LQ) games, [4] extended the scope to infinite-horizon settings, and [hernandez2024ϵ] relaxed the LQ structure via weaker convexity assumptions. Other works addressed different uncertainty sources: [10] proposed adaptive control to mitigate measurement noise, and [19] analyzed the effect of cost parameter variations in discrete-time games.
@@
Nonetheless, the sensitivity of Nash equilibrium to player-specific control deviations, arising from implementation imperfections, remains largely unexplored. Such deviations, e.g., due to actuator errors or delays, are common in engineering systems and can directly degrade performance. Moreover, the discrete-time setting has received little attention in this context. This work analytically investigates how such deviations influence system trajectories and player costs.
@@
To address these effects, we construct a compensation law for the influenced player that uses the nominal equilibrium structure together with measurable deviation information. The resulting controller is not posed as a new Nash equilibrium of the modified interaction; rather, it is a structured compensation policy around the nominal game that mitigates performance loss caused by execution-time deviations. While prior adaptive multi-agent methods exist [9, 22], they do not explicitly target compensation for execution-time control deviations of a specific player, which is addressed in this work.
@@
  • A compensation design for the influenced player, obtained by augmenting the nominal FNE dynamics with measurable deviation dynamics. The resulting policy is shown to be uniquely optimal over a causal affine policy class that includes the uncompensated reference feedback as a special case.
  • A local performance comparison showing that, for sufficiently small deviations, the proposed compensated feedback is no worse than the uncompensated equilibrium-derived feedback to second order, together with a representative numerical example illustrating the resulting performance recovery.
@@
Notation. $\mathbb{R}$ and $\mathbb{Z}$ denote the sets of real and integer numbers, respectively. For a matrix $A$, $A^{\top}$ denotes its transpose; for a square matrix $B$, $B\succ 0$ ($B\succeq 0$) means $B$ is positive definite (positive semi-definite), and $B^{-1}$ denotes its inverse. For $i\in N_{p}=\{1,2\}$, the subscript -i refers to $N_{p}\setminus\{i\}$, i.e., all indices except $i$. The superscript denotes the exact optimal solution. A tilde, as in $\tilde{P}_{i,k}$, denotes auxiliary Riccati-type matrices used to distinguish the coupled finite-horizon game recursion from standard Riccati matrices.
@@ II Preliminaries
We focus on finite-horizon linear quadratic (LQ) games [8] with discrete-time dynamics
@@
We use Nash equilibrium to mean a strategy profile $U_{i}$, where no player can reduce its cost by a unilateral deviation, i.e. $J_{i}(x_{0},U_{i}^{\star},U_{-i}^{\star})\leq J_{i}(x_{0},U_{i},U_{-i}^{\star})$ [12].
@@
The FNE gains $\{K^{\star}_{i,k}\}$ are obtained by a backward Riccati recursion coupled across players [16]. Let $\tilde{P}_{i,N}=Q_{i,N}$ and define, for each $k\in\{0,\dots,N-1\}$ (computed backward),
@@
Under the standard assumptions $Q_{i},Q_{i,N}\succeq 0$, $R_{i}\succ 0$ and nonsingularity of the gain system in Definition 1 at each $k$, the backward sweep $k=N-1\to 0$ yields a unique FNE [8, Sec. 6].
@@
Under conditions 1-3, a unique finite-horizon Nash equilibrium exists. We assume these conditions henceforth [11].
@@ III First-Order Sensitivity of Player 1 Cost under Opponent Deviations
as stated in Definition 1. Substituting (5) into the dynamics (1) yields the time-varying closed-loop matrix
@@
with the convention $\Phi(j,j):=I$.
The state-transition matrix satisfies the standard identities
@@ IV Control Policy Design with Disturbance Compensation
In practice, the control implementation is rarely exact. Differences between the commanded and implemented inputs typically arise from (i) actuator/servo dynamics and inner-loop tracking (first-order lags, rate/slew limits), (ii) computation/communication latency and sample-and-hold effects, and (iii) command filtering/smoothing or safety layers that reshape inputs. These effects induce an input disturbance $w_{k}\;:=\;B_{2}\big(u_{2,k}-u_{2,k}^{\star}\big)$, acting precisely through the channel identified by the sensitivity result.
In practice, $w_{k}$ need not be directly measured at the actuator. Instead, Player 1 may form an estimate $\hat{w}_{k}$ from available information. If the commanded input $u^{\star}_{2,k}=-K^{\star}_{2,k}x_{k}$ is known from the nominal FNE model and the implemented input $u_{2,k}$ can be communicated, sensed, or reconstructed from actuator or vehicle telemetry, then one may compute $\hat{w}_{k}=B_{2}\bigl(u_{2,k}-u^{\star}_{2,k}\bigr)$. When $u_{2,k}$ is not directly available, $\hat{w}_{k}$ can instead be obtained using a standard disturbance observer or state estimator built from the plant model and measured state trajectory. In that case, the proposed CF law is applied using $\hat{w}_{k}$ in place of $w_{k}$, while the present analysis corresponds to the ideal full-information case.
We therefore model $w_{k}$ using simple, identifiable dynamics and equip Player 1 with a Compensated Feedback (CF) law containing a feedforward term $L_{k}w_{k}$ to offset the predictable component of the disturbance. In this way, the feedback term $K_{k}$ regulates the residual error rather than reacting only after the state has already deviated. We first consider a physically grounded actuator-lag model and then narrow it down to a slow-varying error model.
@@ IV-A Compensated Feedback under Lagged Execution Error Dynamics
Inspired by the Disturbance-Based State Compensation (DBSC) approach [15], which is commonly used in traditional state-feedback control design, and given Assumption 1, we define the augmented state as $z_{k}:=[x_{k}^{\top}\ w_{k}^{\top}]^{\top}$.
@@
By (21) in Theorem 1, the gains $\{K_{k},L_{k}\}$ are optimal over the causal affine class (16), which includes the REF baseline $u_{1,k}^{\text{REF}}$ as the special case $L_{k}\equiv 0$. Consequently, when execution-time deviations by Player 2 are measurable, Player 1 can improve upon the uncompensated reference feedback by adopting the compensated law within this policy class. This does not redefine the Nash equilibrium of the underlying game: the FNE is associated with the nominal game without implementation deviations, whereas the proposed law is a compensation strategy for the realized disturbed interaction. The approach is practically relevant because such deviations routinely arise from disturbances, physical limits, and execution imperfections.
@@ IV-B Specialization: Slowly-Varying Error Dynamics
which generally differ from the FNE dynamics $x_{k+1}=\big(A-B_{1}K^{\star}_{1,k}-B_{2}K^{\star}_{2,k}\big)\,x_{k}$. Stability and performance must therefore be assessed for the Linear Time-Varying (LTV) matrix sequence produced by the augmented Riccati recursion.
@@ V Numerical Example: Two Carts Coupled by a Spring–Damper
Let $x:=\begin{bmatrix}p_{1}&v_{1}&p_{2}&v_{2}\end{bmatrix}^{\top}$, where $p_{1}$ and $p_{2}$ are the cart positions, and $v_{1}$ and $v_{2}$ are the corresponding velocities, and $u_{1}$ and $u_{2}$ denote the longitudinal actuation forces applied by Player 1 and 2, respectively. The continuous-time system dynamics are $\dot{x}=A_{c}x+B_{1,c}u_{1}+B_{2,c}u_{2}$ with the matrices:
@@
We compare three implementations over the same horizon and initial condition. Case I (FNE): both players apply their finite-horizon feedback Nash laws computed from the coupled Riccati recursion; no actuator lag is present, and Player 2’s intended command is applied directly. Case II (REF): Player 2’s intended command is subjected to a first-order discrete lag with factor $\alpha$, but Player 1 continues to use its nominal feedback policy; this isolates the performance loss due to the execution lag. Case III (CF): the same actuator lag is present on Player 2, while Player 1 designs a compensated feedback by solving a finite-horizon Riccati recursion on the exact augmented realization of the lagged plant, yielding time-varying gains that act on the state and on the lag-induced mismatch signal. Figure 1 summarizes the two-cart configuration and qualitatively contrasts the resulting final positions under the three cases.
@@
Under the compensated feedback (Case III, dotted), Player 1 modifies its input relative to the nominal FNE action to proactively counteract the actuation shortfall of Player 2. The resulting trajectories stay appreciably closer to the FNE benchmark, especially for Player 1, confirming that CF recovers a substantial fraction of the lag-induced performance loss. Player 2’s cost, however, rises further because it neither alters its strategy nor removes the actuator lag, highlighting the asymmetric benefit in a general-sum game.
Numerically, Table II shows that Player 1’s cost $J_{1}$ increases by $51.36\%$ from Case I to Case II, when lag is present without compensation. Under CF in Case III, Player 1’s cost decreases to $439.67$, yielding a $10.71\%$ improvement relative to the uncompensated lagged case and recovering about $31.56\%$ of the nominal-to-lag performance loss. At the same time, Player 2’s cost $J_{2}$ increases from $492.11$ in Case II to $556.40$ in Case III.
Table II: Finite-horizon performance under the three cases ($\tau=0.8$ s, $\alpha\approx 0.9814$). CF (Case III) reduces Player 1’s lag penalty relative to Case II.
@@
$\mathrm{ISE}_{1}$ 0.1655 0.4566 0.2681
$\mathrm{ISE}_{2}$ 0.1689 0.5076 0.3113
To further assess transient regulation quality, Table II also reports the finite-horizon integral squared error (ISE) for each player, where $\mathrm{ISE}_{1}$ corresponds to Player 1 and $\mathrm{ISE}_{2}$ corresponds to Player 2. These values show the same qualitative improvement for Player 1: relative to Case II, $\mathrm{ISE}_{1}$ is reduced by $41.28\%$ under CF. Interestingly, Player 2’s $\mathrm{ISE}_{2}$ also decreases, by $38.67\%$, even though Player 2’s cost increases. This indicates that the increase in $J_{2}$ is not due solely to poorer regulation. Rather, because Player 2’s objective penalizes both state deviation and control effort, the modified interaction can reduce regulation error while still increasing the overall quadratic cost.
However, the total cost is increased. This tradeoff is expected in the present formulation. The proposed CF law is obtained by minimizing Player 1’s cost over the causal affine class (16) while Player 2 remains fixed at its nominal FNE strategy and still suffers from actuator lag. Therefore, the design is intentionally player-specific rather than socially optimal, and a reduction in $J_{1}$ does not in general imply a reduction in $J_{1}+J_{2}$. Mitigating this efficiency tradeoff would require a joint redesign of both players under implementation dynamics, or alternatively a cooperative/team objective, which lies beyond the scope of the present work.
@@ VI Conclusion
In this paper, we analyzed the impact of control implementation errors in finite-horizon, two-player, nonzero-sum LQ games. We quantified how deviations from one player’s nominal feedback Nash equilibrium input perturb the closed-loop trajectory and affect the other player’s cost. In particular, we derived explicit first-order sensitivity expressions that characterize this dependence in the discrete-time setting.
Motivated by this sensitivity analysis, we constructed a compensated feedback law for the influenced player by augmenting the nominal dynamics with measurable deviation dynamics. The resulting controller is optimal over a causal affine policy class and includes the uncompensated equilibrium-derived feedback as a special case.
A representative numerical example illustrated how the proposed compensation can recover a substantial portion of the performance loss induced by actuator lag. These results highlight the value of exploiting measurable implementation deviations when the exact execution of nominal Nash policies is not possible.
@@
A promising avenue for future research is the development of online estimation techniques combined with adaptive control laws that can anticipate deviations of the other player from the Nash strategy and dynamically adjust to these errors in real time. Furthermore, examining the effects of more complex deviation patterns, such as time-varying, correlated, or structured errors, could offer deeper insights into the robustness and adaptability of game-theoretic control strategies.